OpenStudy (anonymous):

[f*g][2] f[x=2x+1; g[x]=3x-2 how is this set up? no final answer necessary

jhonyy9 (jhonyy9):

first multiplie f with g and after you check what will add this equation for x=2

OpenStudy (anonymous):

you mean if f(x)=2x+1 g(x)=3x-2 fog(2=?)

jhonyy9 (jhonyy9):

is it ok ?

OpenStudy (anonymous):

O_o I'm so lost lol. Is it suppose to look like this, 2(2)+1*3(2)-2??

jhonyy9 (jhonyy9):

tinystar are you there ?

OpenStudy (anonymous):

yeah im here

jhonyy9 (jhonyy9):

this is ok ? you have understood it ?

OpenStudy (anonymous):

no i don't understand it =[

OpenStudy (anonymous):

(f*g)(2) f(x)=2x+1; g(x)=3x-2

jhonyy9 (jhonyy9):

first you need multiplie (2x+1)(3x-2)

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

ok?

jhonyy9 (jhonyy9):

comen

OpenStudy (anonymous):

6x^2-x-2 okay i think i got it. and then substitute in 2 =] THANK YOU

jhonyy9 (jhonyy9):

my pleasure

jhonyy9 (jhonyy9):

this is all ?

OpenStudy (anonymous):

yes